# What makes a good feedforward computational graph?

### v1

@@ Appendix C Adjacency matrix gallery
Appendix D Models and Training
We test graphs on two different tasks: bit set parity, which is well known to be hard for transformer architectures without Chain Of Thought (Kim & Suzuki, 2024) and finding the category of the maximum or second maximum element in the set. We are interested in length generalisation properties of the tasks. So while we train on length up to $256$ elements, we test sequences of up to $1,024$ elements.
D.1 Maximum retrieval
@@
D.2 Parity
We use a model similar to Gemma 2 (Team et al., 2024), but with only only 8 layers and using standard multi-head attention with 8 heads, the vocabulary size was 2, as there are only 2 symbols in a bitstring, and the embedding dimension was $256$. The loss is a crossentropy and is computed over all positions in the sequence, so the model must predict the parity for all prefixes of the input bitstring. To train the model we used the LaProp optimizer (Ziyin et al., 2021) with weight decay and RMSClip (Shazeer & Stern, 2018) for 1 million steps, with batch size of $128$ sequences. Our hyperparameters were:

### v2

@@ Appendix D Adjacency Matrix Gallery
Appendix E Models and Training
We test graphs on two different tasks: bit set parity, which is well known to be hard for transformer architectures without Chain Of Thought (Kim & Suzuki, 2024) and finding the category of the maximum or second maximum element in the set. We are interested in length generalisation properties of the tasks. So while we train on length up to $256$ elements, we test sequences of up to $1,024$ elements. We tuned the order-of-magnitude of the learning rate and the weight decay coefficient using the fully-connected graph experiments only, and then reused the tuned parameters everywhere else.
E.1 Maximum Retrieval
@@
E.2 Parity
We use a model similar to Gemma 2 (Team et al., 2024b), but with only only 8 layers and using standard multi-head attention with 8 heads, the vocabulary size was 2, as there are only 2 symbols in a bitstring, and the embedding dimension was $256$. The loss is a crossentropy and is computed over all positions in the sequence, so the model must predict the parity for all prefixes of the input bitstring. To train the model we used the LaProp optimizer (Ziyin et al., 2021) with weight decay and RMSClip (Shazeer & Stern, 2018) for 1 million steps, with batch size of $128$ sequences. Our hyperparameters were:
@@ E.2 Parity
As an indication of the utility of various feedforward graphs in the setting where nodes correspond to natural language tokens, we also fine-tuned Gemma 2B (Team et al., 2024a) – utilising these graphs as attention masks across all Transformer layers – on the standard Wikipedia dataset [^4] https://www.tensorflow.org/datasets/catalog/wikipedia containing texts obtained from Wikipedia database dumps.
After $3,000$ batches of training, we found that the perplexities obtained by the FS graph were comparable with the fully connected graph, while significantly improving on the line graph—a trend which is consistent throughout training.